Find the intensity of an electric field at a distance of 2× 10^-8 cm from a monovalent ion.The ion is point - charged.


1
Expert's answer
2022-01-31T07:35:18-0500

Solution

Distance d ="2*10^{-8}m"

Monovalent charge q=charge of electrons

So

the intensity of an electric field at this distance

"E=\frac{Kq}{d^2}\\=\frac{9*10^9*1.6*10^{-19}}{(2*10^{-8})^2}"

"=3.6*10^6N\/C."